OpenStudy (anonymous):

Is x + 3 a factor of 3x^3 + 5x^2 + 2x - 12? Please explain! :)

Directrix (directrix):

If the value x = -3 causes 3x^3 + 5x^2 + 2x - 12 to equal 0, then (x+3) is a factor of that expression.

Directrix (directrix):

Evaluate it and see if you get zero.

Directrix (directrix):

You can go here and read about the relationship between zeros and factors of a polynomial (equation). http://www.sosmath.com/algebra/factor/fac02/fac02.html

OpenStudy (anonymous):

I still don't understand! :( Is this what I was supposed to do? 3(-3)^3 + 5(-3)^2 + 2(-3) - 12 = -9^3 - 15^2 - 6 - 12 = -729 - 225 - 6 - 12 = -972

OpenStudy (anonymous):

 You have a remainder of 30 over your original statement, so no it is not a factor.